# On the robustness of modeling grounded word learning through a child’s egocentric input

### v1

@@ 4.5 Finding 5: Robust evidence of learnability from training models across different children’s datasets
Figure 6: Labeled-S Classification Performance between different combinations of Whisper transcriptions, across three different model configurations. The left-most bar in each group represents performance from Figure 5. The two additional bars in each group explored modifications to the language encoder, either replacing the Embedding encoder with a Transformer (T) or with a Transformer and an additional language modeling head (T+LM). Models were trained with three random seeds and results were averaged.
@@
Finally, to test whether our findings generalize beyond a single child’s linguistic environment, we examined model performance across multiple children’s datasets. This addresses one of the key limitations from our previous analyses, which focused exclusively on learning from just one child’s data. While it has been sufficient to demonstrate learnability from relatively generic learning mechanisms, it fails to address how generalizable and robust these findings are. Here, we leverage Whisper’s ability to generate automatic transcriptions for all babies from the SAYCam dataset: S, A and Y. In this analysis, we train models on either S-Whisper, A-Whisper, or Y-Whisper.
@@
In order to properly evaluate these models trained on other children’s data, we also required a separate set of evaluation trials that were drawn from each child’s videos respectively. This set of evaluations also consists of 4-way classification trials, but generated from a larger set of 41 categories, see Appendix C for additional dataset creation details. Notably, these evaluations were designed in a similar fashion to the S-Whisper-Disjoint split described above: the evaluation frames were generated from the “test” splits of each child’s data, such that they were independent from anything seen during training or validation. [^4] However, due to differences in dataset splitting, the the new set of evaluation frames derived from baby S are not completely independent from training on S-Whisper, but they are for the A-Whisper and Y-Whisper datasets. These evaluation datasets all contain the same set of 41 categories that were present in all three children’s visual and linguistic environments, enabling both a more comprehensive and comparable method of comparison for grounded word learning across children. Furthermore, we generate a different evaluation dataset from each child’s test frames and only evaluate models trained on that child’s data to their corresponding evaluation set, which we call Labeled-S-V2, Labeled-A and Labeled-Y respectively.
@@
Results for these evaluations are shown in Figure 7. At the highest level, we find that all model configurations and dataset splits, when evaluated on their corresponding evaluation split, achieves a classification accuracy higher than random chance (25%). Overall, these results show that word-referent mappings are robustly learnable with the use of multimodal neural networks, not just in the case of one child’s dataset (Vong et al.,, 2024), but instead across multiple independent children’s datasets and across a variety of different model configurations. Looking more closely at the performance across dataset splits, we still see some qualitative differences in performance. Models trained on data from baby S and baby Y achieve higher levels of accuracy (at 51% and 49% respectively), while the best model configuration for baby A only achieves 43% accuracy. Similar to the previous set of results in comparing model variants on baby S’s dataset splits, here we also find that no specific model configuration achieves superior classification performance, but instead performance is generally in the same range for each split. Overall, these findings provide strong evidence that multimodal neural network based approaches can be used to acquire word-referent mappings, both for a larger set of categories than previously explored, and across models trained from different children’s datasets.
@@
Figure 7: Classification accuracy across models trained from Whisper transcriptions of each child in SAYCam. Each model was evaluated on a new set containing 41 different categories, shared across the three evaluation datasets but using video frames derived from each child’s videos (see Figure 1 for some example evaluation frames). Models were trained with three random seeds and results were averaged.
@@ 5 Discussion
In this work, we explored the robustness of simulated word learning from developmentally plausible datasets using multimodal neural networks with minimal inductive biases. We used automatic speech transcription methods to transcribe the entire SAYCam dataset, increasing the number of available transcribed utterances by a factor of 7, and use these automatically transcribed utterances to generate a number of multimodal training and evaluation datasets. We tested 3 different model configurations varying the type of language encoder and loss functions, against 5 different dataset splits across 3 babies. Overall, our results showed that all model configurations, across all three datasets, demonstrated that word-referent mappings were learnable using multimodal neural networks. Our findings build upon the results from Vong et al., (2024) — although here we used almost twice as many visual concepts — highlighting that the kind and quantity of visual and linguistic information present in children’s environments is indeed sufficient to acquire and generalize many word-referent mappings, without the need for additional linguistic inductive biases. Across multiple dataset splits, we show that the learned word embeddings generalize to visual instances to entirely new videos not seen during the training process.
While the use of automatic transcriptions for multimodal modeling has previously been explored (Roy and Pentland,, 2002; Harwath et al.,, 2018), it has seen less uptake in the field of language acquisition, likely due to some of the difficulties in obtaining accurate and representative transcripts (although see Long et al., (2024) for another recent example). In this work, we find that training multimodal models from automatic transcripts enabled the training of separate models for each of the three children’s dataset from the SAYCam corpora, that were previously inaccessible due to the efforts required for manual speech transcription. Furthermore, by comparing models trained from the same subset of videos from baby S either from manual vs. automated transcriptions, we find only a minor gap in classification performance. However, one advantage in our particular case for examining word-referent mappings is that the visual input of children changes relatively slowly given the egocentric nature of these datasets. Thus, Whisper’s imperfect utterance timings, combined with the wide range from which we sample paired video frames, is apparently sufficient for learning.
@@
While our results showed that the specific model configuration used for each dataset resulted in minor variations in classification performance, we found that no specific model configuration was dominant across all datasets. However, in the second set of evaluations, comparing performance across models trained across the three children with the same set of evaluation categories (although with different frames), we observed that performance on models trained on baby A were much worse compared to babies S or Y. One hypothesis we arrived at after watching a number of the videos is that there are far fewer instances of visual-linguistic alignment present in the training set. In particular, baby A’s recordings (and transcribed utterances) often involved both parents having a discussion that was unrelated to the captured visual input, providing limited instances for multi-modal alignment for the model. While prior research has shown that children can acquire new words from overhearing, they often still required the novel words to be relevant in the current visual context (Akhtar,, 2005; Gampe et al.,, 2012).
@@
Overall, our results provide a roadmap for studying multimodal language acquisition in developing children. Compared to Vong et al., (2024), we explore datasets from multiple children and numerous model types and configurations, showing robustness to the acquisition of word-referent mappings across all of these configurations. Nevertheless, there still remain many other interesting research opportunities at the intersection of developmentally rich datasets and modern machine learning methods to further close the data-efficiency gap and understand how humans effectively learn language from so little. In particular, we hope future work will build on our learnability results to develop more psychologically and developmentally plausible cognitive models (Portelance and Jasbi,, 2024), incorporating more realistic constraints as well as mechanisms for learning from actions and social interaction.
@@ Acknowledgments
We are grateful to the authors of the SAYCam article (Sullivan et al.,, 2022), and the volunteers who contributed to the data set. Thanks to Emin Orhan for training separate vision transformers from each child’s data, which were used as the vision backbone for our multimodal neural networks. Additional thanks to Qiwen Zhang, Hailey He, Manli Zhao, Luyang Shang and Michael Picheny for their guidance on using Whisper and WhisperX to help automatically transcribe the SAYCam dataset, and their efforts to manually verify the quality of the transcriptions for this work.
@@ Appendix A Generating Evaluation Datasets
In this work, we generated three new evaluation datasets (Labeled-S-V2, Labeled-A and Labeled-Y) to evaluate our multimodal models, which was performed using a mixture of both automatic and manual filtering. To accomplish this, we created new labeled datasets derived from the corresponding video frames from (primarily) the test set videos for each child, [^5] The test set videos were unused during the training procedure, although for baby S there was some partial overlap for baby S creating three evaluation datasets consisting of 41 shared categories with 100 evaluation trials each, for a total of 4100 tirals.
These evaluation sets were created by first generating a set of target words or concepts to filter for. We generated this list by first considering the set of single word concepts from the MacArthur-Bates Communicative Development Inventory (MCDI), followed by filtering to (1) only consider words whose concreteness score was higher than 4.8 (using concreteness scores from Brysbaert et al., (2014)), (2) present in all three children’s vocabularies, and (3) whose part-of-speech tag was either a NOUN or VERB, producing a set of 156 potential target labeled concepts to scan for.
Next, we used the largest available CLIP model (ViT-L/14@336px) (Radford et al.,, 2021), and computed a similarity score (via cosine similarity) between every frame from each child’s set of test set frames and the set of 156 target concepts (using the text embedding of the target label directly), providing us with an automated method to detect potential labeled frames. Using a similarity threshold of 0.24 (based on qualitative observations), we selected the highest scoring label for any given video frame above this threshold, which provided us with a set of labeled frames for each child. Using this as a starting point, we then performed a second round of manual filtering, by examining every automatically labeled instance to determine whether it was valid or not for the corresponding category label. After this manual filtering step, we performed a final filtering step to only include categories for which there were 10 or more instances in each of the three children’s labeled datasets, leading to 41 categories in total.
@@ Appendix B Generating the CLIP-filtered Validation Sets
Due to the limited signal from the noisy frame-utterance pairs that were automatically extracted, we needed a stronger signal to determine whether models were learning during the training process. To facilitate this, we leveraged CLIP to separately create filtered versions of the validation set for each data split, which was performed once as a pre-processing step prior to model training, ensuring that only frame-utterance pairs whose similarity exceeded a threshold were used to compute the validation loss. We used CLIP ViT-B/32 (Radford et al.,, 2021), calculating the cosine similarity between a randomly selected frame and its associated utterance. We filtered the validation set to only include frame-utterance pairs whose similarity scores were greater than 0.24, [^6] This threshold was determined via manual inspection from a small number of similarity scores. thereby removing noisy or uninformative pairs for model validation during the training process. A comparison table for the size of the CLIP-filtered validation sets for each dataset split can be found in Table 2.
Appendix C Evaluation Dataset Categories
Labeled-S Categories (22 categories)
ball, basket, car, cat, chair, computer, crib, door, floor, foot, ground, hand, kitchen, paper, puzzle, road, room, sand, stairs, table, toy, window
Labeled-S-V2, Labeled-A and Labeled-Y Categories (41 categories)
arm, ball, banana, bed, bedroom, blanket, book, boots, bottle, bowl, box, bread, bucket, cereal, cow, cup, door, finger, foot, hair, horse, jeans, kitchen, knee, leg, monkey, napkin, oven, paper, pen, pillow, porch, shirt, shoe, shoulder, sock, sun, table, towel, toy, window
Appendix D Tables

### v2

@@ 4.5 Finding 5: Robust evidence of learnability from training models across different children’s datasets
Figure 6: Labeled-S Classification Performance between different combinations of Whisper transcriptions, across three different model configurations. The left-most bar in each group represents performance from Figure 5. The two additional bars in each group explored modifications to the language encoder, either replacing the Embedding encoder with a Transformer (T) or with a Transformer and an additional language modeling head (T+LM). Models were trained with three random seeds. Error bars show bootstrapped 95% confidence intervals over category-level accuracies.
@@
To test whether our findings generalize beyond a single child’s linguistic environment, we examined model performance across multiple children’s datasets. This addresses one of the key limitations from our previous analyses, which focused exclusively on learning from just one child’s data. While it has been sufficient to demonstrate learnability from relatively generic learning mechanisms, it fails to address how generalizable and robust these findings are. Here, we leverage Whisper’s ability to generate automatic transcriptions for all babies from the SAYCam dataset: S, A and Y. In this analysis, we train models on either S-Whisper, A-Whisper, or Y-Whisper.
@@
In order to properly evaluate these models trained on other children’s data, we also required a separate set of evaluation trials that were drawn from each child’s videos respectively. This set of evaluations also consists of 4-way classification trials, but generated from a larger set of 41 categories, see Appendix D for additional dataset creation details. Notably, these evaluations were designed in a similar fashion to the S-Whisper-Disjoint split described above: the evaluation frames were generated from the “test” splits of each child’s data, such that they were independent from anything seen during training or validation, enabling the measurement of within-child generalization. [^7] However, due to differences in dataset splitting, the new set of evaluation frames derived from baby S are not completely independent from training on S-Whisper, but they are for the A-Whisper and Y-Whisper datasets. These evaluation datasets all contain the same set of 41 categories that were present in all three children’s visual and linguistic environments, enabling both a more comprehensive and comparable method of comparison for grounded word learning across children. Furthermore, we generate a different evaluation dataset from each child’s test frames and only evaluate models trained on that child’s data to their corresponding evaluation set, which we call Labeled-S-V2, Labeled-A and Labeled-Y respectively.
@@
Results for these evaluations are shown in Figure 7. At the highest level, we find that all model configurations and dataset splits, when evaluated on their corresponding evaluation split, achieve a classification accuracy higher than random chance (25%). Overall, these results show that word-referent mappings are robustly learnable with the use of multimodal neural networks, not just in the case of one child’s dataset (Vong et al., 2024), but instead across multiple independent children’s datasets and across a variety of different model configurations. Looking more closely at the performance across dataset splits, we still see some qualitative differences in performance. Models trained on data from baby S and baby Y achieve higher levels of accuracy (at 51% and 49% respectively), while the best model configuration for baby A only achieves 43% accuracy. Furthermore, and consistent with the previous set of results in comparing model variants, here we also find that no specific model configuration achieves superior classification performance.
@@
Figure 7: Classification accuracy across models trained from Whisper transcriptions of each child in SAYCam. Each model was evaluated on a new set containing 41 different categories, shared across the three evaluation datasets but using video frames derived from each child’s videos (see Figure 1 for some example evaluation frames). Models were trained with three random seeds. Error bars show bootstrapped 95% confidence intervals over category-level accuracies.
What drives the variation in classification performance across models trained on different children’s datasets? A qualitative inspection of baby A’s videos revealed that a large number of training examples contained parental discussions that were unrelated to the current visual context, providing noisy and irrelevant signals for grounded multimodal learning. To quantify the degree of vision-language alignment present in each child’s training dataset, we used CLIP ViT-B/32 to compute the distribution of cosine similarities between training frames and their paired utterances. While mean CLIP similarities were nearly identical across children (baby S: 0.222, baby A: 0.220, baby Y: 0.216), the count of high-quality aligned examples (CLIP similarity >0.24>0.24) varied substantially. Baby S’s training data contained 9,320 highly-aligned pairs, baby A’s training data had 3,223 pairs, and baby Y’s training data had 3,514 pairs, corresponding to 7.60%, 4.43% and 4.94% of their respective training datasets. The relative ranking of the number and proportion of high-quality training examples matches our observed classification performance, suggesting that the absolute number of high-quality examples, rather than the average quality, is the key driver behind successful multimodal word learning.
4.6 Finding 6: Learned word-referent mappings also transfer across children and to novel image domains
Finally, we examine generalization of word-referent mappings more broadly. First, we examine transfer to other children’s visual environments (cross-child generalization), and second, to an entirely different image domain (out-of-distribution generalization).
Figure 8: Cross-child classification accuracy across models. This figure presents classification accuracy scores when models trained on one child’s data are evaluated on another child’s evaluation set. Rows indicate the training dataset, while columns indicate the evaluation dataset. Diagonal entries represent the degree of within-child generalization, while off-diagonal entries represent the degree of cross-child generalization. All models show some degree of cross-child generalization, although some more than others.
To examine whether models can generalize across children’s visual environments, we tested models on each of the three evaluation datasets from the previous section, regardless of the dataset each model was trained on. That is, models trained on baby S’s videos were evaluated on Labeled-A and Labeled-Y, et cetera. Results are shown in Figure 8. The values along the diagonals represent the classification accuracy results from the previous section, corresponding to models trained and evaluated on the same child. However, the values on the off-diagonals represent the classification accuracy of cases where models were trained and evaluated on different children’s data. Examining these results, we see a broadly similar pattern, while cross-child generalization shows reduced performance compared to within-child generalization. However, the resulting averaged classification accuracy scores remained well above the 25% chance level, indicating that these learned word-referent mappings do indeed transfer over to other children’s visual environments to some degree. Furthermore, and consistent with the results from the previous section, we also see better cross-child generalization performance for models trained on babies S and Y, compared to models trained on baby A.
Figure 9: Out-of-distribution generalization to the Konkle Objects dataset. Each model was evaluated on the same set of evaluation trials containing novel images on white backgrounds, across 60 different categories. Models trained on baby S show partial out-of-distribution performance, while models trained on babies A and Y are modestly above chance. Error bars show bootstrapped 95% confidence intervals over category-level accuracies.
In an even stricter test of out-of-distribution generalization, we also tested models on an adapted version of the same Konkle Objects dataset used in Vong et al. (2024). All models were tested on the same set of evaluation trials, consisting of 60 categories (present in all three children’s vocabularies) of naturalistic objects on white backgrounds across 1,784 trials, none of which were seen during training. Results are shown in Figure 9. While models trained on data from baby S showed some limited generalization around 36 - 37% classification accuracy, models trained on A and Y showed more limited generalization at 30% and 28% average classification accuracy respectively, modestly above chance-level. Consistent with our other evaluations, model architecture had minimal impact on performance, with training data characteristics remaining the primary factor driving performance differences. Together, these results show that all models can generalize beyond their training data, but that performance also progressively decreases from within-child to cross-child to out-of-distribution generalization, reflecting the increasing domain shift at each step.
@@ 5 Discussion
In this work, we explored the robustness of simulated word learning from developmentally plausible datasets using multimodal neural networks with minimal inductive biases. We used automatic speech transcription methods to transcribe the entire SAYCam dataset, increasing the number of available transcribed utterances by a factor of 7, and use these automatically transcribed utterances to generate a number of multimodal training and evaluation datasets. We tested 3 different model configurations varying the type of language encoder and loss functions, against 5 different dataset splits across 3 babies. Overall, our results showed that all model configurations, across all three datasets, demonstrated that word-referent mappings were learnable using multimodal neural networks. Our findings build upon the results from Vong et al. (2024) — although here we used a broader set of evaluation datasets, some with twice as many concepts — highlighting that the kind and quantity of visual and linguistic information present in children’s environments is indeed sufficient to acquire and generalize many word-referent mappings, without the need for additional linguistic inductive biases. Across multiple dataset splits, we show that the learned word embeddings generalize to visual instances across videos, across children and across image domains. Models successfully classify held-out frames from the same child, transfer to the visual environments of other children, and show modest above-chance performance on novel, naturalistic object stimuli, demonstrating some degree of transfer throughout.
@@
While our results showed that the specific model configuration used for each dataset resulted in minor variations in classification performance, we found that no specific model configuration was dominant across all datasets. However, in the second set of evaluations, comparing performance across models trained across the three children with the same set of evaluation categories (although with different frames), we observed that performance on models trained on baby A were much worse compared to babies S or Y. One hypothesis we arrived at after watching a number of the videos is that there are far fewer instances of visual-linguistic alignment present in the training set. In particular, baby A’s recordings (and transcribed utterances) often involved both parents having a discussion that was unrelated to the captured visual input, providing limited instances for multi-modal alignment for the model. While prior research has shown that children can acquire new words from overhearing, they often still required the novel words to be relevant in the current visual context (Akhtar, 2005; Gampe et al., 2012).
While the use of automatic transcriptions for multimodal modeling has previously been explored (Roy and Pentland, 2002; Harwath et al., 2018), it has seen less uptake in the field of language acquisition, likely due to some of the difficulties in obtaining accurate and representative transcripts (although see Long et al. (2024) for another recent example). In this work, we find that training multimodal models from automatic transcripts enabled the training of separate models for each of the three children’s dataset from the SAYCam corpora, that were previously inaccessible due to the efforts required for manual speech transcription. Furthermore, by comparing models trained from the same subset of videos from baby S either from manual vs. automated transcriptions, we find only a minor gap in classification performance. However, one advantage in our particular case for examining word-referent mappings is that the visual input of children changes relatively slowly given the egocentric nature of these datasets. Another advantage is that our evaluations focus on common object categories with sufficient exemplars during training, masking some of the issues with errors during transcription, or dealing with more uncommon words. Thus, Whisper’s imperfect utterance timings, combined with the wide range from which we sample paired video frames, turns out to suffice for word learning. Finally, we note that our approach and current findings are specific to English, and generalizing this approach to other languages may prove to be challenging, given the high variability in accuracy of automated speech transcriptions across different languages, especially low-resource ones (Radford et al., 2023).
@@
Overall, our results provide a roadmap for studying multimodal language acquisition in developing children. Compared to Vong et al. (2024), we explore datasets from multiple children and numerous model types and configurations, showing robustness to the acquisition of word-referent mappings across all of these configurations. Nevertheless, there still remain many other interesting research opportunities at the intersection of developmentally rich datasets and modern machine learning methods to further close the data-efficiency gap and understand how humans effectively learn language from so little. In particular, we hope future work will build on our learnability results to develop more psychologically and developmentally plausible cognitive models (Portelance and Jasbi, 2024), incorporating more realistic attention and memory constraints, mechanisms for learning from actions and social interaction, and additional endogenous sources of linguistic input (Zhao et al., 2025). Another natural extension is learning from raw speech rather than text transcriptions, which would allow models to leverage phonological and prosodic cues. While modeling from speech introduces additional challenges due to acoustic variability (Merkx et al., 2019), recent progress in speech-image learning suggests this is an increasingly tractable research direction (Harwath et al., 2016; Chrupała, 2022).
@@ Acknowledgments
We are grateful to the authors of the SAYCam article (Sullivan et al., 2022), and the volunteers who contributed to the data set. Thanks to Emin Orhan for training separate vision transformers from each child’s data, which were used as the vision backbone for our multimodal neural networks. Additional thanks to Qiwen Zhang, Hailey He, Manli Zhao, Luyang Shang and Michael Picheny for their guidance on using Whisper and WhisperX to help automatically transcribe the SAYCam dataset, and their efforts to manually verify the quality of the transcriptions for this work.
@@ Appendix A Generating Evaluation Datasets
A.1 Labeled-S-V2, Labeled-A and Labeled-Y
In this work, we generated three new evaluation datasets (Labeled-S-V2, Labeled-A and Labeled-Y) to evaluate our multimodal models, which was performed using a mixture of both automatic and manual filtering. To accomplish this, we created new labeled datasets derived from the corresponding video frames from (primarily) the test set videos for each child, [^8] The test set videos were unused during the training procedure, although for baby S there was some partial overlap. creating three evaluation datasets consisting of 41 shared categories with 100 evaluation trials each, for a total of 4100 trials.
These evaluation sets were created by first generating a set of target words or concepts to filter for. We generated this list by first considering the set of single word concepts from the MacArthur-Bates Communicative Development Inventory (MCDI), followed by filtering to (1) words whose concreteness score was higher than 4.8 (using concreteness scores from Brysbaert et al. (2014)), (2) words present in all three children’s vocabularies, and (3) words whose part-of-speech tag was either a NOUN or VERB, producing a set of 156 potential target labeled concepts to scan for.
Next, we used the largest available CLIP model (ViT-L/14@336px) (Radford et al., 2021), and computed a similarity score (via cosine similarity) between every frame from each child’s test set frames and the set of 156 target concepts (using the text embedding of the target label directly), providing us with an automated method to detect potential labeled frames. Using a similarity threshold of 0.24 (based on qualitative observations), we selected the highest scoring label for any given video frame above this threshold, which provided us with a set of labeled frames for each child. Using this as a starting point, we then performed a second round of manual filtering, by examining every automatically labeled instance to determine whether it was valid or not for the corresponding category label. After this manual filtering step, we performed a final filtering step to only include categories for which there were 10 or more instances in each of the three children’s labeled datasets, leading to 41 categories in total.
@@ A.1 Labeled-S-V2, Labeled-A and Labeled-Y
A.2 Konkle Objects
The Konkle Objects evaluation dataset was derived from a naturalistic image dataset containing multiple exemplars from 200 common object categories on white backgrounds from Konkle et al. (2010). Following Vong et al. (2024), which used a 64-category subset from this dataset for evaluating models trained on baby S, we created an adapted version that was suitable for comparing across children to measure out-of-distribution generalization. Specifically, we selected the subset of object categories that were present in the vocabulary of all three children, resulting in 60 object categories. This choice of common categories ensured that models trained on any of the three children’s datasets could be evaluated on this dataset.
For each of the exemplars in the 60 object categories, we generated 5 independent evaluation trials, consisting of one target image along with three additional foil images randomly selected from the other categories. This evaluation set consists of 1,784 evaluation trials across the 60 categories.
@@ Appendix B Generating the CLIP-filtered Validation Sets
Due to the limited signal from the noisy frame-utterance pairs that were automatically extracted, we needed a stronger signal to determine whether models were learning during the training process. To facilitate this, we leveraged CLIP to separately create filtered versions of the validation set for each data split, which was performed once as a pre-processing step prior to model training, ensuring that only frame-utterance pairs whose similarity exceeded a threshold were used to compute the validation loss. We used CLIP ViT-B/32 (Radford et al., 2021), calculating the cosine similarity between a randomly selected frame and its associated utterance. We filtered the validation set to only include frame-utterance pairs whose similarity scores were greater than 0.24, [^9] This threshold was determined via manual inspection from a small number of similarity scores. thereby removing noisy or uninformative pairs for model validation during the training process. A comparison table for the size of the CLIP-filtered validation sets for each dataset split can be found in Table 2.
Appendix C Word Similarity Evaluation Results
In addition to our classification-based evaluations, we also conducted a separate set of word similarity evaluations to examine whether the use of a Transformer-based architecture, or language modeling objective produces richer semantic representations. Following the methodology of Zhuang et al. (2024b), we used the word similarity benchmark MTest-3000 (Bruni et al., 2012), containing human similarity judgments for 3000 word pairs. We filtered this set down to only include pairs of words where both words had an age of acquisition (AoA) of 10 or below (Kuperman et al., 2012), and were present in the vocabulary of all three children, resulting in 725 word similarity pairs for evaluation.
Figure 10: Word similarity evaluation across models and datasets. Spearman correlation between model word embedding similarities and human similarity judgments, using the MTest-3000 word similarity benchmark filtered to word pairs with AoA $\leq$ 10 and present in each model’s vocabulary. Joint models (CVCL+T+LM) consistently show higher correlations than contrastive-only models across all three children’s datasets.
For each word pair, we computed the cosine similarity between the embeddings of the language encoder of each model, and measured the Spearman correlation between model similarities and human similarity judgments. Results are shown in Figure 10. Across each child, we see that the joint models (CVCL+T+LM) showed stronger correlations with human similarity judgments, compared to either of the contrastive-only models (CVCL and CVCL+T), indicating that the additional language modeling objective captures some additional semantic structure that contrastive-only models do not.
Appendix D Evaluation Dataset Categories
Labeled-S Categories (22 categories): ball, basket, car, cat, chair, computer, crib, door, floor, foot, ground, hand, kitchen, paper, puzzle, road, room, sand, stairs, table, toy, window
Labeled-S-V2, Labeled-A and Labeled-Y Categories (41 categories): arm, ball, banana, bed, bedroom, blanket, book, boots, bottle, bowl, box, bread, bucket, cereal, cow, cup, door, finger, foot, hair, horse, jeans, kitchen, knee, leg, monkey, napkin, oven, paper, pen, pillow, porch, shirt, shoe, shoulder, sock, sun, table, towel, toy, window
Konkle Objects Categories (60 categories): airplane, apple, bagel, ball, balloon, basket, bed, bell, bike, bill, bird, boot, bottle, bowl, bucket, butterfly, button, cake, camera, cat, chair, cheese, clock, cookie, crib, dog, doll, fan, guitar, hat, jacket, juice, key, knife, leaves, lock, meat, necklace, pants, pen, phone, pipe, pizza, ring, rock, rug, sandwich, shoe, socks, spoon, stool, tape, tent, train, tree, trumpet, turtle, tv, umbrella, watch
Appendix E Tables